# SpecServe: Efficient and SLO-Aware Large Language Model Serving with Adaptive Speculative Decoding

### v1

@@ 5 Adaptive Speculative Decoding
Based on the relationship between efficiency and speculative length (§4.3), we propose adaptive speculative decoding. To implement this framework, we developed three key components. The first is the Adaptive Drafter (§5.1), which incorporates efficiency estimation into the drafting phase, achieving forward-looking step-level speculative length control. The second is the Confidence Prior Verifier (§5.2), which prioritizes the verification of tokens with high acceptance rates, achieving fine-grained request-level speculative length control. Lastly, the SLO-Aware Efficiency Estimator (§5.3) evaluates the efficiency of speculative decoding and achieves SLO awareness.
5.1 Adaptive Drafter
@@
0:  Draft model $M_{d}$, Request set $reqs$, Confidence score of previous steps $prev\_conf$
0:  Draft token set $dtss$, Acceptance rate set $arss$
1:  $best\_thpt\leftarrow-1$, $dtss\leftarrow\{\}$, $arss\leftarrow\{\}$
2:  $prev\_confs\leftarrow\text{list}(prev\_conf)\times\text{lens}(reqs)$
3:  $arss.\text{append}(prev\_confs)$
@@
6:     if $thpt>best\_thpt$ then
@@
To address this challenge, we adopted a "predict-execute-correct" strategy. As demonstrated in Alg. 1, we first use the average confidence scores from historical data to estimate the confidence scores that the next step of the draft model might produce, and based on this, predict the efficiency of speculative decoding. If the estimation indicates that continuing the execution of the draft model would improve the efficiency, the speculative controller executes the model, taking into consideration the newly generated draft tokens and their confidence scores, while also updating the collection of acceptance rates and the efficiency. If the estimation suggests that it is not appropriate to continue execution, the drafting phase is stopped. Through this approach, we are able to enable the draft model to rapidly generate draft tokens that are more likely to be accepted.
@@
0:  Target model $M_{t}$, Request set $reqs$, Draft token set $dtss$, Acceptance rate set $arss$
0:  Output token set $otss$
1:  $best\_thpt\leftarrow\text{throughput\_estimation}(reqs,arss)$
@@
6:     if $thpt>best\_thpt$ then
@@
13:  $otss\leftarrow M_{t}(reqs,dtss)$
5.2 Confidence Prior Verifier
We developed a module named the confidence prior verifier, which introduces an innovative request-level speculative length control strategy by decoupling the drafting and verification phases of speculative decoding. Currently, speculative length control is typically implemented on step-level [22], which limits the ability to adjust speculative lengths for different requests within the same batch individually. The confidence prior verifier achieves request-level speculative length control indirectly by prioritizing the verification of draft tokens anticipated to have a higher acceptance rate, thereby validating different numbers of draft tokens for different requests. Considering that the verification phase may occupy up to 90% of the total time for speculative decoding, this differentiated verification strategy among requests efficiently realizes fine-grained speculative length control.
@@ 5.2 Confidence Prior Verifier
5.3 SLO-Aware Efficiency Estimator
@@
0:  Request set $reqs$, Acceptance rate set $arss$, SLO’s TPOT constraint $tpot$
0:  Estimated goodput $gdpt$
1:  $t\leftarrow\text{estimate\_time}(reqs)$
@@
Drawing from the model of accepted token numbers presented in §4.2, the SLO-aware efficiency estimator utilizes the acceptance rate predicted by confidence to determine the number of tokens to be accepted at the current step. Subsequently, in conjunction with the speculative decoding time model outlined in §4.1, it calculates the time required for speculative decoding based on the current workload. Through these two parameters, the estimator is able to estimate the speculative decoding throughput at the current step, thereby assisting SpecServe in achieving the optimal efficiency of speculative decoding.
Furthermore, we guide SpecServe in selecting the appropriate speculative length according to SLO requirements by setting constraints. Considering that in a batch of requests, some may only generate a single accepted token, making their TPOT equivalent to the time of speculative decoding. Therefore, to ensure that the time for each step of speculative decoding does not exceed the TPOT constraint specified by the SLO, strict control is necessary. If the time for speculative decoding exceeds the SLO limit, the throughput will be set to -1 to prevent the adaptive drafter from proceeding. In this manner, SpecServe can adjust the speculative length based on SLO constraints to better meet the SLO.
6 Implementation
SpecServe is implemented by extending the vLLM [17] with approximately 2000 lines of Python code. vLLM is a well-known production-level serving system that efficiently manages memory through PagedAttention and integrates online inference techniques such as continuous batching. At the speculative decoding mode, the lookahead scheduler within the vLLM divides the memory into two equal parts, allocating one for the KV Cache of the draft model and the other for the KV Cache of the target model. When a new batch of requests comes in, the adaptive drafter performs the draft model autoregressively through an "predict-execute-correct" process, maintaining a collection of acceptance rates until generating a set of draft tokens with the optimal efficiency of speculative decoding at the step level. Following this, the confidence prior verifier eliminates some draft tokens with low acceptance rates that hinder the efficiency of speculative decoding, and then inputs the set of draft tokens with the optimal efficiency at the request level into the target model for verification. After generating the results of the speculative decoding for this step, these results are integrated into the input of the speculative decoding for the next step.
@@ 6 Implementation
7 Evaluation
7.1 Experimental Setup
Table 1: Model pairs and hardware.
Target Model Draft Model Hardware
Vicuna-7B-v1.5 Vicuna-68M 1×L40
1×A800
Vicuna-33B-v1.3 Vicuna-160M 2×A800
4×A800
Llama-3.1-70B-Instruct Llama-3.2-1B-Instruct 8×A800
@@ 7.1 Experimental Setup
Model pairs and computing platforms. We conducted tests on three model pairs, which were completed across five different computational platforms. Specifically, we selected three target models: Vicuna-7B-v1.5 [43], Vicuna-33b-v1.3, and Llama-3.1-70B-Instruct [9], and matched them with Vicuna-68M [36], Vicuna-160M, and Llama-3.2-1B-Instruct as their respective draft models. Considering the memory resources, the model pair for the 7B target model operated on a single L40 or A800 GPU, the model pair for the 33B target model operates on either 2 $\times$ A800 GPUs or 4 $\times$ A800 GPUs, while the model pair for the 70B target model runs on 8 $\times$ A800 GPUs. In all these setups, the draft model operates with tensor parallelism set to 1, and only the target model is shared across multiple GPUs. We provide detailed specifications of the setup in Table 1.
@@
Workloads. To evaluate the performance of SpecServe in real-world online inference systems, we selected three real-world production traces from the AzureLLMInferenceDataset2023 [29], specifically from the code and conversation categories, as well as mooncake [30], to simulate different request arrival patterns. To maintain system load within a normal range, we chose 10-minute segments with moderate request rates and adjusted the request rate of mooncake to one-third of its original rate. Consequently, these three traces respectively simulated scenarios of bursty requests (trace 1), high request rates (trace 2), and low request rates (trace 3).
To simulate the diversity of request types in real online inference systems, we created a mixed request set based on SpecBench [35], incorporating a variety of request types such as translation, summarization, question answering, mathematical reasoning, and retrieval-augmented generation. This request set was composed of 480 randomly selected instances from six widely used databases (MT-bench [43], WMT14 DE-EN, CNN/Daily Mail [27], Natural Questions [16], GSM8K [7], and DPR [14]).
@@
Figure 9: Speedup in end-to-end latency across different traces, model pairs and computing platforms.
@@
Figure 10: SLO attainment and speedup in end-to-end latency across different traces, model pairs and computing platforms.
@@
Baselines. We compared SpecServe with four baseline methods: (1) autoregressive decoding; (2) vanilla speculative decoding with preset speculative lengths of 1, 3, and 5; (3) threshold-based dynamic speculative decoding, a dynamic speculative decoding method that has been integrated into the transformer library [34], representing the state-of-the-practice method; (4) SpecInfer [26], a state-of-the-art speculative decoding method integrated into inference systems. Due to the inefficiency of tree-based inference strategies (with a tree width of 3, the acceptance rate of draft tokens does not exceed 18% [26]), we set the tree width to 1 and name it SpecInfer*.
Metrics. To evaluate the reduction in latency, we measure the end-to-end latency speedup relative to autoregressive decoding. Drawing upon prior research [44], we established the TTFT at 200ms and the TPOT at 30ms as our SLO metrics. By adjusting the scale factor within the range of 0.8 to 1.4, we simulated varying degrees of SLO stringency.
7.2 Overall Performance of SpecServe
@@
As illustrated in Fig. 9-11, we first evaluated the end-to-end latency speedup and SLO attainment of all methods under real-world production traces, and assessed the end-to-end latency speedup of all methods across various batch sizes.
Consistent low inference latency. As shown in Fig. 9, we observe that SpecServe’s speedup ranged from 1.1× to 1.8× relative to autoregressive decoding across 15 experimental setups, demonstrating its stability in reducing latency. In contrast, vanilla speculative decoding methods (with speculative lengths of 1, 3, 5) have a speedup ratio of less than 1 in 7%, 13%, and 40% of the experimental setups, respectively, indicating that maintaining low latency becomes increasingly challenging with longer preset lengths. Threshold-based dynamic method does not achieve the expected outcome in 40% of the setups, suggesting that a single confidence score is insufficient for accurately determining the optimal speculative length. SpecInfer* is the least stable, primarily due to its inefficient speculative strategy. By adapting to different request patterns and system configurations, SpecServe consistently optimized end-to-end latency, making it a reliable inference engine suitable for real-world production environments.
Compared with the optimal methods in various scenarios, SpecServe outperforms the baseline method in terms of latency in 10 out of 15 experimental setups, and matches the latency performance of the baseline method in the remaining 5 setups. From a system configuration perspective, SpecServe demonstrates a more significant latency reduction in resource-constrained scenarios, such as the deployment of a 33B+160M model pair on a 2×A800 computing platform compared to the same model pair on a 4×A800 computing platform. From the perspective of workload, when facing scenarios with bursty requests (trace 1), SpecServe is capable of achieving an additional speedup of 5% to 10% compared to other baseline methods. This indicates that SpecServe’s fine-grained speculative length control method can adaptively determine the optimal speculative length, especially suitable for resource-limited and complex request pattern production environments.
Figure 11: Speedup in end-to-end latency across various batch sizes.
High SLO attainment. Fig. 10 demonstrates the SLO attainment and speedup of different methods under request burst scenarios (trace 1) with varying SLO constraints. As depicted, SpecServe’s SLO attainment rate exceeds that of other methods by 5% to 38% across all SLO constraints, and its speedup relative to autoregressive decoding surpasses other methods by 10% to 61%. Furthermore, SpecServe can adjust its speculative length according to different SLO constraints, thereby altering its speedup. For instance, under loose SLO constraints, SpecServe can achieve an additional 2% speedup. This indicates that SpecServe not only attains a higher speedup while maintaining or exceeding the SLO attainment of autoregressive decoding but also flexibly adjusts according to varying SLO constraints to optimize performance. This characteristic effectively mitigates the issue of speculative decoding, where pursuing a higher speedup often results in a lower SLO attainment, making SpecServe a fast and SLO-aware solution.
Batch adaptivity. As shown in Fig. 11, under arbitrary batch sizes composed of different requests, SpecServe achieves the highest speedup among all compared baseline methods. Its superior batch adaptability accounts for the demonstrated stable latency reduction and low-latency characteristics. In contrast, vanilla speculative decoding with a preset length cannot match the batch adaptability of dynamic methods. These static method either employ conservative speculative lengths to maintain a relatively stable reduction in latency or use aggressive speculative lengths to gamble for a higher speedup, unable to achieve both simultaneously. Obviously, dynamic methods also require a reasonable speculative length control scheme. Threshold-based approaches can only offer random improvements in speedup and are incapable of maintaining stable latency reduction. In summary, SpecServe’s superior batch adaptability is particularly crucial for production environments, as it ensures that SpecServe can maintain efficient batch processing capabilities most of the time, even in the face of significant fluctuations in batch size.
7.3 Ablation Study
@@
Table 2: The variations in average speculative length, average number of accepted draft tokens, average draft token acceptance rate, and speedup across different traces for various dynamic methods.
@@
Threshold-based 6.54 1.49 0.20 0.54×
Trace 1 SpecServe- 1.64 0.63 0.40 1.09×
SpecServe 1.59 0.63 0.42 1.13×
Threshold-based 5.67 1.44 0.20 0.85×
Trace 2 SpecServe- 3.26 1.18 0.39 1.34×
SpecServe 3.17 1.16 0.40 1.38×
Threshold-based 3.84 1.34 0.31 1.51×
Trace 3 SpecServe- 4.20 1.34 0.33 1.57×
SpecServe 4.08 1.33 0.34 1.58×

### v2

@@ 5. Adaptive Speculative Decoding
Based on the relationship between efficiency and speculative length (§4.3), we propose adaptive speculative decoding. To implement this framework, we developed three components. The first is the Adaptive Drafter (§5.1), which incorporates efficiency estimation into the drafting phase, achieving forward-looking step-level speculative length control. The second is the Confidence Prior Verifier (§5.2), which prioritizes the verification of tokens with high acceptance rates, achieving fine-grained request-level speculative length control. Lastly, the SLO-Aware Efficiency Estimator (§5.3) evaluates the efficiency of speculative decoding and achieves SLO awareness.
5.1. Adaptive Drafter
@@
0: Draft model $M_{d}$, Request set reqsreqs, Confidence score of previous steps prev_confprev\_conf
0: Draft token set dtssdtss, Acceptance rate set arssarss
$1best\_thpt\leftarrow$-1, $dtss\leftarrow\{\}$, $arss\leftarrow\{\}$
[…]
@@
6:  if thpt>best_thptthpt>best\_thpt then
@@
To address this challenge, we adopted a ”predict-execute-correct” strategy. As demonstrated in Alg. 1, we first use the average confidence scores from historical data to estimate the confidence scores that the next step of the draft model might produce, and based on this, predict the efficiency of speculative decoding. If the estimation indicates that continuing the execution of the draft model would improve the efficiency, the speculative controller executes the model, taking into consideration the newly generated draft tokens and their confidence scores, while also updating the collection of acceptance rates and the efficiency. If the estimation suggests that it is not appropriate to continue execution, the drafting phase is stopped. This approach enables the draft model to rapidly generate draft tokens that are more likely to be accepted.
@@
0: Target model $M_{t}$, Request set reqsreqs, Draft token set dtssdtss, Acceptance rate set arssarss
0: Output token set otssotss
[…]
@@
6:  if thpt>best_thptthpt>best\_thpt then
@@
[…]
5.2. Confidence Prior Verifier
We developed a module named the confidence prior verifier, which introduces an innovative request-level speculative length control strategy by decoupling the drafting and verification phases of speculative decoding. Currently, speculative length control is typically implemented on step-level (Liu et al., 2024b), which limits the ability to adjust speculative lengths for different requests within the same batch individually. The confidence prior verifier achieves request-level speculative length control indirectly by prioritizing the verification of draft tokens anticipated to have a higher acceptance rate, thereby validating different numbers of draft tokens for different requests. Considering that the verification phase may occupy up to 90% of the decoding time, this differentiated verification strategy among requests efficiently realizes fine-grained speculative length control.
@@ 5.2. Confidence Prior Verifier
5.3. SLO-Aware Efficiency Estimator
@@
0: Request set reqsreqs, Acceptance rate set arssarss, SLO’s TPOT constraint tpottpot
0: Estimated goodput gdptgdpt
[…]
@@
Drawing from the model of accepted token numbers presented in §4.2, the SLO-aware efficiency estimator utilizes the acceptance rate predicted by confidence to determine the number of tokens to be accepted at the current step. In conjunction with the speculative decoding time model outlined in §4.1, it calculates the time required for speculative decoding based on the current workload. Through these two parameters, the estimator can estimate the speculative decoding throughput at the current step, thereby assisting AdaSpec in achieving the optimal efficiency of speculative decoding.
Furthermore, we guide AdaSpec in selecting the appropriate speculative length according to SLO requirements by setting constraints. Considering that in a batch of requests, some may only generate a single accepted token, making their TPOT equivalent to the time of speculative decoding. Therefore, to ensure that the time for each step of speculative decoding does not exceed the TPOT constraint specified by the SLO, strict control is necessary. If the time for speculative decoding exceeds the SLO limit, the throughput will be set to -1 to prevent the adaptive drafter from proceeding. In this manner, AdaSpec can adjust the speculative length based on SLO constraints to better meet the SLO.
6. Implementation
AdaSpec is implemented by extending the vLLM (Kwon et al., 2023) with approximately 2000 lines of Python code. vLLM is a well-known production-level serving system that efficiently manages memory through PagedAttention and integrates online inference techniques such as continuous batching. At the speculative decoding mode, the lookahead scheduler within the vLLM divides the memory into two equal parts, allocating one for the KV Cache of the draft model and the other for the KV Cache of the target model. When a new batch of requests comes in, the adaptive drafter performs the draft model autoregressively through an ”predict-execute-correct” process, maintaining a collection of acceptance rates until generating a set of draft tokens with the optimal efficiency of speculative decoding at the step level. Following this, the confidence prior verifier eliminates some draft tokens with low acceptance rates that hinder the efficiency of speculative decoding, and then inputs the set of draft tokens with the optimal efficiency at the request level into the target model for verification. After generating the results of the speculative decoding for this step, these results are integrated into the input of the speculative decoding for the next step.
@@ 6. Implementation
7. Evaluation
7.1. Experimental Setup
Table 1. Model pairs and hardware.
Target Model (TP) Draft Model (TP) Hardware (VRAM)
Vicuna-7B-v1.5 (1) Vicuna-68M (1) 1×L40 (48GB)
1×A100 (40GB)
Vicuna-33B-v1.3 (4) Vicuna-160M (1) 4×A100 (160GB)
Llama-3.3-70B (8) Llama-3.2-1B (1) 8×A100 (320GB)
@@ 7.1. Experimental Setup
Model pairs and computing platforms. We conducted tests on three model pairs, which were completed across four different computational platforms. Specifically, we selected three target models: Vicuna-7B-v1.5 (Zheng et al., 2023), Vicuna-33b-v1.3, and Llama-3.3-70B-Instruct (Dubey et al., 2024), and matched them with Vicuna-68M (Yang et al., 2024), Vicuna-160M, and Llama-3.2-1B-Instruct as their respective draft models. Considering the memory resources, the model pair for the 7B target model operated on a single L40 or A100 (40GB) GPU, the model pair for the 33B target model operates on 4 $\times$ A100 (40GB) GPUs, while the model pair for the 70B target model runs on 8 $\times$ A100 (40GB) GPUs. In all these setups, the draft model operates with tensor parallelism set to 1, and only the target model is shared across multiple GPUs. We provide detailed specifications in Tab. 1.
@@
Workloads. To evaluate the performance of AdaSpec in real-world cloud inference systems, we selected three representative production traces from BurstGPT (Wang et al., 2024) and Mooncake (Qin et al., 2025) to simulate request arrival patterns. To ensure the system load remains within normal operating ranges, we extracted 10-minute windows with moderate request rates for evaluation. As shown in Fig. 1, these three traces simulate the following scenarios: (1) significant request rate variations (trace 1), (2) periodic request rate fluctuations (trace 2), and (3) rapid request rate oscillations (trace 3).
To accurately simulate the diversity of requests in online inference scenarios, we developed a mixed request set based on SpecBench (Xia et al., 2024) encompassing six distinct task categories: translation (WMT14 DE-EN), summarization (CNN/Daily Mail (Nallapati et al., 2016)), question answering (Natural Questions (Kwiatkowski et al., 2019)), mathematical reasoning (GSM8K (Cobbe et al., 2021)), retrieval-augmented generation (DPR (Karpukhin et al., 2020)), and general dialogue (MTbench (Zheng et al., 2023)). For each task category, we randomly sampled 80 instances from its corresponding dataset, totaling 480 request instances. Since the number of requests varies across traces, we sampled instances from this mixed set with equal probability to simulate real-world request patterns. Given that different target models employ distinct tokenizers and output modes, we present only the input and output length distributions corresponding to the Vicuna-7B-v1.5 model, as shown in Fig. 9.
Figure 9. Input and output length distributions corresponding to the Vicuna-7B-v1.5 model, showcasing the workload’s key patterns.
@@
Figure 10. Speedup in end-to-end latency across different traces, model pairs and computing platforms.
@@
Figure 11. SLO attainment and speedup in end-to-end latency across different traces, model pairs and computing platforms.
@@
Baselines. We compared AdaSpec with four baseline methods: (1) autoregressive decoding; (2) vanilla speculative decoding with preset speculative lengths of 1, 3, and 5; (3) threshold-based dynamic speculative decoding, a dynamic speculative decoding method that has been integrated into the transformer library (Wolf et al., 2020), representing the state-of-the-practice method; (4) SpecInfer (Miao et al., 2024), a state-of-the-art speculative decoding method integrated into inference systems. Due to the inefficiency of tree-based inference strategies (with a tree width of 3, the acceptance rate of draft tokens does not exceed 18% (Miao et al., 2024)), we set the tree width to 1 and name it SpecInfer*.
Metrics. To measure the latency reduction effect, we use the speedup of end-to-end latency relative to autoregressive decoding as our primary metric. Since our focus is on the benefits of speculative decoding in improving TPOT, we set the SLO for TPOT as the P90 TPOT value of autoregressive decoding, employing the SLO attainment rate to assess service reliability (Zhong et al., 2024; Liu et al., 2024a). Furthermore, to evaluate the robustness of our approach under varying SLO constraints, we adjust the scale factor within a range of 0.8 to 1.2, simulating different levels of SLO strictness.
7.2. Overall Performance of AdaSpec
@@
As illustrated in Fig. 10-12, we first evaluated the end-to-end latency speedup and SLO attainment of all methods under real-world production traces, and assessed the end-to-end latency speedup of all methods across various batch sizes.
Consistent low inference latency. As shown in Fig. 10, AdaSpec demonstrates significant advantages across all 12 experimental configurations. Its core innovation lies in the fine-grained dynamic speculative length control mechanism that enables stable performance improvements across diverse complex scenarios. Compared to autoregressive decoding, AdaSpec achieves 1.23× to 1.79× speedup, proving the universal superiority of its speculative decoding framework. Against baseline methods with fixed speculative lengths, AdaSpec delivers an additional 7%-18% speedup through real-time speculative optimization, highlighting the critical value of its dynamic adjustment mechanism. The 16%-66% additional speedup over SpecInfer* reveals the inefficiency of verifying large batches of draft tokens in complex request patterns. Notably, AdaSpec achieves 9% to 58% greater acceleration than threshold-based dynamic methods, demonstrating that relying solely on confidence scores cannot accurately determine optimal speculative lengths, thus validating the advancement of AdaSpec’s efficiency-optimized speculative decisions. By adapting to various request patterns and system configurations, AdaSpec consistently optimizes end-to-end latency, establishing itself as a reliable inference engine for real-world cloud production environments.
From a workload adaptability perspective, AdaSpec exhibits unique elastic capabilities: its 18%-33% speedup advantage in scenarios with drastic request rate fluctuations (trace 1) demonstrates exceptional adaptability to bursty loads; while maintaining stable 13%-17% speedup in rapidly changing environments (trace 3), confirming the robustness of its dynamic decision mechanism. Even in scenarios with pronounced periodic patterns (trace 2), AdaSpec breaks through the performance ceiling of baseline methods to achieve additional 10%-14% speedup, attributable to its fine-grained speculative length control strategy. Comparative analysis across these three scenarios shows AdaSpec effectively adapts to varying system loads, with performance benefits amplifying as load variability increases - a characteristic that makes AdaSpec particularly suitable for complex production environments.
In terms of system configurations, AdaSpec excels across hardware environments with varying arithmetic intensity: from 20% performance improvement in 1×L40 configurations to consistent 15% acceleration in 1×A100 setups. Its adaptive speculative strategy effectively overcomes the generalization limitations of traditional methods across hardware with different computational characteristics, making it an ideal choice for production deployment.
Figure 12. Speedup in end-to-end latency across various batch sizes.
@@
Table 2. The variations in average speculative length, average number of accepted draft tokens, average draft token acceptance rate, and speedup across different traces for various dynamic methods.
@@
Threshold-based 5.97 1.41 0.24 1.16×
Trace 1 AdaSpec- 3.57 1.22 0.34 1.71×
AdaSpec 3.42 1.20 0.35 1.73×
Threshold-based 6.56 1.42 0.22 1.22×
Trace 2 AdaSpec- 3.13 1.17 0.37 1.40×
AdaSpec 2.96 1.14 0.39 1.42×
Threshold-based 5.26 1.38 0.26 1.66×
Trace 3 AdaSpec- 3.73 1.26 0.34 1.78×
AdaSpec 3.63 1.25 0.35 1.79×
High SLO attainment. Fig. 11 presents the performance of various methods under a scenario with significant request rate fluctuations (trace 1) and varying SLO constraints. The experimental results demonstrate that AdaSpec exhibits significant advantages in achieving high SLO attainment. When the SLO constraint is set to the P90 TPOT value of autoregressive decoding (SLO scale=1), AdaSpec is the only method achieving an SLO satisfaction rate exceeding 90%, with a speedup ratio over autoregressive decoding that outperforms other methods by 6% to 66%. Notably, under all SLO constraints, AdaSpec consistently achieves higher SLO satisfaction rates (up to 38% improvement) while also delivering the most substantial speedup gains (up to 66% improvement).
Furthermore, AdaSpec demonstrates unique adaptive capabilities, dynamically adjusting its speculative length based on different SLO constraints to optimize performance. For instance, under the most lenient SLO constraint (SLO scale=1.2), AdaSpec achieves an additional 2% speedup compared to the strictest condition (SLO scale=0.8). This characteristic allows it to successfully address a critical challenge in speculative decoding—where pursuing higher speedup ratios typically leads to degraded SLO satisfaction rates. By employing SLO-aware adaptive speculative length control, AdaSpec not only maintains or surpasses the SLO satisfaction levels of autoregressive decoding but also achieves superior acceleration performance. As a result, it emerges as an ideal solution that balances both efficiency and SLO compliance.
Batch adaptivity. In dynamic production environments, fluctuations in batch size significantly impact system performance. As illustrated in Fig. 12, AdaSpec demonstrates exceptional batch adaptability when handling arbitrary batch sizes with varying request compositions, achieving the highest speedup ratio among all baseline methods. This superior adaptability is reflected in the system’s stable latency reduction and low-latency characteristics.
Compared to dynamic approaches, traditional fixed-length speculative decoding methods exhibit notable deficiencies in batch adaptability. These static methods typically face a dilemma: employing a conservative speculative length may maintain relatively stable latency reduction but fails to achieve optimal speedup, while adopting an aggressive speculative length may yield higher speedup at the cost of latency stability. Notably, even dynamic methods require an appropriate speculative length control mechanism to achieve optimal performance. For instance, threshold-based dynamic methods can provide sporadic speedup improvements but struggle to maintain consistent latency reduction.
AdaSpec overcomes these limitations through its innovative dynamic adjustment mechanism. Its exceptional batch adaptability ensures efficient batch processing capability even under drastic batch size fluctuations—a critical feature for real-world production environments. This enables AdaSpec to deliver consistently stable and high-performance operation across diverse complex scenarios.
7.3. Ablation Study